OpenStudy (anonymous):

Solve the equation. Then check your solution. 3.7(-83x-3)=3.7(7x-30) a. 0.7 c. 1.3 b. –0.3 d. 0.3

OpenStudy (anonymous):

Just expand both sides and solve. 3.7(-83x-3)=3.7(7x-30) -307.1x-11.1 = 25.9x-111

OpenStudy (anonymous):

Can you solve from there?

OpenStudy (anonymous):

can you go just on step further but witout the answer for me or no

OpenStudy (anonymous):

i need to be able to put it in a calculator

OpenStudy (jjcoolja12):

Let's solve your equation step-by-step. 3.7(−83x−3)=3.7(7x−30) Step 1: Simplify both sides of the equation. 3.7(−83x−3)=3.7(7x−30) Simplify: (Show steps) −307.1x−11.1=25.9x−111 Step 2: Subtract 25.9x from both sides. −307.1x−11.1−25.9x=25.9x−111−25.9x −333x−11.1=−111 Step 3: Add 11.1 to both sides. −333x−11.1+11.1=−111+11.1 −333x=−99.9

OpenStudy (jjcoolja12):

solve the rest of it

OpenStudy (anonymous):

Put like terms on the same side. -307.1x-11.1 = 25.9x-111 333x=99.9

OpenStudy (jjcoolja12):

that's -33x=99.9 not postive 333x

OpenStudy (jjcoolja12):

*-333x

OpenStudy (anonymous):

No I added -307.1x.

OpenStudy (anonymous):

And I added 111.

OpenStudy (jjcoolja12):

Let's solve your equation step-by-step. 3.7(−83x−3)=3.7(7x−30) Step 1: Simplify both sides of the equation. 3.7(−83x−3)=3.7(7x−30) Simplify: (Show steps) −307.1x−11.1=25.9x−111 Step 2: Subtract 25.9x from both sides. −307.1x−11.1−25.9x=25.9x−111−25.9x −333x−11.1=−111 Step 3: Add 11.1 to both sides. −333x−11.1+11.1=−111+11.1 −333x=−99.9 Step 4: Divide both sides by -333. −333x −333 = −99.9 −333

OpenStudy (anonymous):

When you divide 99.9 by 333 (or -99.9 by -333) what do you get?

OpenStudy (jjcoolja12):

what you get

OpenStudy (anonymous):

-0.3

OpenStudy (anonymous):

or 0.3

OpenStudy (jjcoolja12):

good

OpenStudy (jjcoolja12):

Let's solve your equation step-by-step. 3.7(−83x−3)=3.7(7x−30) Step 1: Simplify both sides of the equation. 3.7(−83x−3)=3.7(7x−30) Simplify: (Hide steps) (3.7)(−83x)+(3.7)(−3)=(3.7)(7x)+(3.7)(−30)(Distribute) −307.1x+−11.1=25.9x+−111 −307.1x−11.1=25.9x−111 Step 2: Subtract 25.9x from both sides. −307.1x−11.1−25.9x=25.9x−111−25.9x −333x−11.1=−111 Step 3: Add 11.1 to both sides. −333x−11.1+11.1=−111+11.1 −333x=−99.9 Step 4: Divide both sides by -333. −333x −333 = −99.9 −333 x=0.3 Answer: x=0.3

OpenStudy (jjcoolja12):

https://www.youtube.com/watch?v=YZBStgZGyDY https://www.youtube.com/watch?t=92&v=f15zA0PhSek this can help you in the future

OpenStudy (anonymous):

thanks both of you

OpenStudy (jjcoolja12):

your welcome

OpenStudy (phi):

notice it's easier if the first step is divide both sides by 3.7

OpenStudy (phi):

you would then get -83x-3=7x-30 which you (with practice) can do in your head. 27= 90 x x= 27/90 divide top and bottom by 9 x= 3/10